### v3

@@ 4.2.1 GENERATIVE PATCH DISTRIBUTION MATCHING (GPDM)
The process of synthetic data generation, as demonstrated in Fig. 2, utilizes one-shot generative models. Corollary 2 establishes a theoretical bound on the distribution distance between the original dataset and the generated dataset. This ensures that while the dataset is expanded using one-shot generative techniques, the deviation from the original’s distribution remains within a controlled limit. In implementing this method, we leverage architectures that are originally designed for single image generation, adapting them to treat the reshaped samples as $45\times 45$ images. This reshaping approach is substantiated by Corollary 2 and is intended to align with the operational needs of the generative model. Following this step, we reverse the reshaping process for the subsequent phase, i.e. the classification task.
In this work, we use the algorithm in [27], namely GPDM, which is based on minimizing the Sliced Wasserstein Distance (SWD) between the patches of real and generated images. The algorithm also utilizes a multi-scale hierarchical strategy to learn the global structure of the image in the early stages of training and its hyperparameter are detailed in Table 2.
@@ 4.2.2 CONDITIONAL DEEP GENERATIVE MODELS
We compare conditional deep generative models trained on clean versus unclean data and augment training data using the mixup method to prevent overfitting, smooth labels, and enhance model generalization. Our examination extends to detailed architectures and strategies for Conditional Generative Adversarial Networks (CGAN) and Conditional Variational Autoencoders (CVAE), focusing on layer structures and hyperparameters.
Data Cleaning
Outlier sensitivity is a critical concern in the training of deep generative models, particularly in low-data regimes where the limited number of training samples can magnify the impact of anomalous data. Such anomalies can lead to mode collapse, a situation where the model fails to capture the diversity of the data distribution and instead focuses on a few modes, significantly degrading the model’s performance. Removing outliers is thus essential to ensure that the training process accurately represents the underlying data distribution, promoting a more robust and generalizable model.
Therefore, we compare the results of conditional deep generative models in two scenarios: using clean and unclean training data. The cleaning process employs the Median Absolute Deviation (MAD) method, using a threshold of 3 MADs to determine and remove outliers. For each class, outliers are identified by calculating the Euclidean norms (L2 norms) of the sample features. A sample is considered an outlier if the absolute deviation of its norm from the median norm exceeds the threshold of three times the MAD. This selective exclusion enhances the robustness and reliability of the models by ensuring that they are trained on the most representative data points.
Augmenting Training Data
In our approach to enhancing model performance, we employ the mixup method  [19] to expand our training dataset to 5000 samples. This process involves mixing pairs of data and label samples from the original dataset based on weights derived from a symmetric Beta distribution, specifically Beta($\alpha=0.2$, $\beta=0.2$).
This augmentation strategy is critical before employing deep generative models, which require substantial amounts of data. By increasing our dataset size through mixup, we address the high data demands of these models and promote label smoothing. This preparation helps mitigate overfitting and enhances the model’s generalization capabilities by exposing it to a more varied set of training examples, which is essential for effectively training deep generative models to learn robust features.
CGAN
CGAN is a type of GAN that conditions its generation process on additional information like labels or data from other modalities. This conditioning leads to more controlled and targeted generation of data, allowing the network to produce specific types of outputs based on the given conditions [30]. We enhance the implementation of CGAN by integrating Wasserstein GAN with Gradient Penalty (WGAN-GP), which improves training stability and mitigates mode collapse, thereby ensuring higher quality and diversity in the generated outputs  [31]. In our architecture of CGAN, the selection of depth and width for layers follows a pyramidical structure, with an increasing or decreasing format to control processing of information, where each layer refines the output of its predecessor. We also employ label injection across all linear layers, not just the initial ones; see Table 3 where $C\in\{4,10\}$ denotes class size. This approach strengthens the labels’ influence throughout the network. This integration of labels into the network’s layers, as evidenced in literature, bolsters the conditioning effect, aligning outputs more closely with specified labels [32, 33, 34, 35].
Our hyperparameter configuration includes the use of the LeakyReLU activation function, set with a slope of 0.2 for negative input values, to prevent the ”dying ReLU” problem as well as to capture non-linearities in the data  [36, 37]. Batch size is determined as 64 and the learning rate for the Adam optimizer is set to 0.0005 for both discriminator and generator.
@@ 7 Tanh - - - - - -
CVAE
CVAE, a variant of Variational Autoencoders, distinctively incorporates conditional data such as labels or contextual inputs to steer its generative mechanism. This incorporation of conditional information is central to CVAE’s functionality, enabling it to generate data that is both precise and contextually relevant, allowing the network to create outputs finely tuned to the specified conditions. Similar to CGAN, our CVAE architecture embodies a tiered layer structure, mirroring a pyramid-like progression in the depth and width of layers with labels incorporated in all linear layers; see Table 4.
CVAE utilizes a dual-component loss function: a reconstruction loss, ensuring the fidelity of the generated data to the original input, and a Kullback-Leibler (KL) divergence loss, which maintains the efficiency and effectiveness of the latent space representation  [38, 18]. To dynamically balance the contribution of the KL divergence loss across training epochs, the CVAE employs a cyclical annealing schedule, with the KL weight peaking at its maximum value of 1.0 every 10 epochs. This sinusoidal modulation facilitates a smoother and more controlled integration of the KL divergence into the overall loss, enhancing the stability and quality of the latent space learning process  [39]. The configurations for our model include using a LeakyReLU activation function with a slope of 0.2, a batch size of 64, and an Adam optimizer with a learning rate of 0.0001.
@@ 4.3 CLASSIFIER
We opt for a pipeline that integrates the Synthetic Minority Over-sampling Technique (SMOTE)  [40, 41] with XGBoost’s Classifier  [42, 43]. This strategic choice targets class imbalance, a pervasive challenge in machine learning, using SMOTE to bolster the representation of minority classes effectively. Coupled with XGBoost, renowned for its efficiency, scalability, and adaptability across various applications, we employ an ensemble method that significantly enhances accuracy and generalization. This pipeline aligns seamlessly with our overarching research priorities, allowing to channel our efforts and resources more effectively into the core aspect of our work - synthetic data generation - while still ensuring reliable classification results.
In our setup, 20% of the training data is reserved for five random validation sets, over which results are averaged to enhance robustness and reliability for every case. We incorporate a patience of 10 evaluations during validation to determine early stopping, further refining our model’s performance. Our model is configured with the “multi:softprob” objective, suitable for multiclass problems as it outputs a probability for each class. It evaluates using the “mlogloss” and “merror” metrics; “mlogloss” measures the accuracy of the probability outputs while “merror” tracks the misclassification rate, together providing a comprehensive view of model performance across multiple classes.
@@ 4.4 EVALUATION METRICS
The evaluation is conducted on ten tasks: 4-classes and 10-classes UAV (and flight mode) identification, each having five cases of generation on 10, 20, 30, 40, and 50% of the complete data. For each case, we use a randomly sampled 5-fold cross-validation [44] setting and report the average scores over all of the folds.
Four metrics are used to evaluate classification using the eight cases of generated data. As we are dealing with a multilabel classification problem, the following scores are calculated for each label, then are averaged by the weight of their support (the number of true instances for each label). Denote the number of true positive, true negative, false positive, and false negative data as $TP$, $TN$, $FP$, and $FN$, respectively. By definition, we have:
$\begin{split}&\text{Accuracy}=(TP+TN)/(TP+TN+FP+FN)\\ &\text{Precision}=TP/(TP+FP)\\ &\text{Recall}=TP/(TP+FN)\\ &F_{1}=2\times(\text{Precision}\times\text{Recall})/(\text{Precision}+\text{Recall})\end{split}$ (30)
@@ 5.1 VALIDATION OF BOUNDED DIVERGENCE
In Section 3, we established an upper bound for the distributional distance between synthetic and real data, focusing on the distribution of patches within samples. Specifically, the bound presented in Corollary 2 ensures that two-dimensional arrays, such as images, generated by one-shot generative models like GPDM, do not exhibit a distributional divergence from real arrays beyond a predefined threshold. In this subsection, we demonstrate that such threshold holds in practice.
Let $W$ represent the distributional distance between synthetic and real datasets, and $\delta$ denote the expected distributional distance between their respective patches. We demonstrate the relationship of $W/\delta$ to its upper bound, as detailed in Eq. (22). Fig. 5 confirms that Eq. (22) is applicable across various scenarios, encompassing different class counts, patch sizes used in generation, proportions of real data utilized, and the patch size $(n^{\prime})$ employed for calculating $\delta$. The figure presents the mean of $W/\delta$, calculated over five folds, and also shows this mean value plus ten times the standard deviation of $W/\delta$.
Despite the consistent applicability of the bound across all scenarios, it may not always appear tight. In this work, the observed phenomenon can primarily be attributed to the multiscale approach of GPDM. This approach initially concentrates on capturing the global structure of samples at a lower scale. This process effectively reduces $W$ to a greater extent. Also, $\delta$ is subject to overestimation using limited patches within each sample, because the Wasserstein distance is sensitive to the tails of distributions. Limited samples may not adequately represent these tails, leading to an overestimation of the actual distributional distance. Additionally, in high-dimensional spaces, the curse of dimensionality can inflate perceived differences, further contributing to overestimation. Sampling variability, especially when capturing more extreme values or outliers, can also skew the perceived distribution, resulting in an overestimated Wasserstein distance between the respective patches $(\delta)$.
Therefore, we have successfully established and validated an upper bound for the distributional distance between synthetic and real data using one-shot generative models, specifically focusing on GPDM. Our investigations confirm the practical applicability of the established bound across a range of scenarios, as demonstrated in the consistency of $W/\delta$’s adherence to its theoretical upper limit. However, it is important to recognize that certain factors, such as the multiscale nature of GPDM and the limited patch counts within each sample, can lead to the underestimation of $W/\delta$’s in empirical circumstance.
@@ 5.2 CLASSIFICATION PERFORMANCE
We can evaluate the identification of UAV types and flight modes by employing synthetic RF transforms across various scenarios. These scenarios are distinguished by the proportion of data used in the synthesis process and the generation methodology. Specifically, we assess classification performance by comparing a complete dataset and its subsets sampled at different proportions with their corresponding synthetically generated counterparts, utilizing a range of models.
The classification performance metrics are presented in Figs. 6 and 7, which both incorporate results from complete and subsampled datasets for comparison. In this context, the complete datasets serve as our ground truth, while the subsampled datasets function as baselines. The metrics for the complete data remain constant since data proportion is not a factor. On the other hand, the metrics for the subsampled data improve with the use of more data.
@@
Figure 6: The performace of UAV identification using synthetic transformed RF signals generated by GPDM.
@@
Figure 7: The performace of UAV identification using synthetic transformed RF signals generated by CGAN and CVAE.
Fig. 6 illustrates the performance metrics of classification using synthetic RF transformations generated by the one-shot GPDM approach. Generally, the performance metrics of the synthetic data fall between those of the baseline and the ground truth, demonstrating the effectiveness of the one-shot generative approach for data synthesis. The impact of patch size on various metrics is significant, especially in the 10-class scenario where utilizing a patch size of 5 is more successful than larger patch sizes in expanding the support of class-wise distributions. The benefit of choosing a smaller patch size becomes less pronounced as the proportion of data accessed increases (approaching 50%), where further extrapolation of distributional support no longer offers additional benefits.
Fig. 7 illustrates the performance metrics of classifications using synthetic RF transformations generated by conditional deep learning models such as CGAN and CVAE. Contrary to the GPDM approach, CGAN and CVAE underperform, not even reaching the baseline metrics, primarily due to their inefficiency in scenarios with limited data. In such low-data regimes, the outcome where deep learning models tend to memorize small training samples represents the best-case scenario. This already precarious situation is further aggravated if the training data is not meticulously clean. The presence of noise and outliers in the data poses significant challenges, making it a struggle for these models to even achieve basic memorization of the training samples, let alone generalize effectively.
One-shot generative data augmentation, in contrast to CGAN and CVAE, excels particularly in environments with outliers. Its robustness arises from its capacity for internal learning, where the model extracts and utilizes the intrinsic characteristics of a single example to generate a broad range of synthetic samples. This approach focuses on the core attributes of the sample, effectively minimizing the influence of outliers. By capturing and extrapolating the essential features, the one-shot method not only improves the diversity and representativeness of the synthetic data but also significantly enhances its utility for training models in scenarios plagued by data anomalies.
@@
The comparative analysis, depicted in Fig. 6 and Fig. 7, reveals significant insights into the performance of the one-shot generative approach using GPDM and conditional deep generators such as CGAN and CVAE. While the one-shot GPDM method consistently demonstrates superior handling and generation of synthetic data in the presence of outliers and limited data scenarios, the performance of CGAN and CVAE lags, especially when data cleanliness and abundance are issues. This study underscores the critical importance of choosing appropriate data synthesis techniques based on the specific challenges and data environment encountered. It suggests that one-shot generative methods, with their robust internal learning mechanisms, are not only more resilient but also more adaptable to diverse and challenging data landscapes, making them a preferable choice in situations where data quality cannot be guaranteed.

### v4

@@ 4.2.1 Generative patch distribution matching (GPDM)
The process of synthetic data generation, as demonstrated in Fig. 2, utilizes one-shot generative models. Corollary 2 establishes a theoretical bound on the distribution distance between the original dataset and the generated dataset. This ensures that while the dataset is expanded using one-shot generative techniques, the deviation from the original’s distribution remains within a controlled limit. In implementing this method, we leverage architectures that are originally designed for single image generation, adapting them to treat the reshaped samples as $45\times 45$ images. This reshaping approach is substantiated by Corollary 2 and is intended to align with the operational needs of the generative model. Following this step, we reverse the reshaping process for the subsequent phase, i.e. the classification task. In this work, we use the algorithm in [28], namely GPDM, which is based on minimizing the SWD between the patches of real and generated images. The algorithm also utilizes a multi-scale hierarchical strategy to learn the global structure of the image in the early stages of training and its hyperparameter are detailed in Table 2.
@@ 4.2.2 Conditional deep generative models
We compare conditional deep generative models trained on clean versus unclean data and augment training data using the mixup method to prevent overfitting, smooth labels, and enhance model generalization. Our examination extends to detailed architectures and strategies for Conditional Generative Adversarial Networks (CGAN) and Conditional Variational Autoencoders (CVAE), focusing on layer structures and hyperparameters. Outlier sensitivity is a critical concern in the training of deep generative models, particularly in low-data regimes where the limited number of training samples can magnify the impact of anomalous data. Such anomalies can lead to mode collapse, a situation where the model fails to capture the diversity of the data distribution and instead focuses on a few modes, significantly degrading the model’s performance. Removing outliers is thus essential to ensure that the training process accurately represents the underlying data distribution, promoting a more robust and generalizable model. Therefore, we compare the results of conditional deep generative models in two scenarios: using clean and unclean training data. The cleaning process employs the Median Absolute Deviation (MAD) method, using a threshold of 3 MADs to determine and remove outliers. For each class, outliers are identified by calculating the Euclidean norms (L2 norms) of the sample features. A sample is considered an outlier if the absolute deviation of its norm from the median norm exceeds the threshold of three times the MAD. This selective exclusion enhances the robustness and reliability of the models by ensuring that they are trained on the most representative data points. In our approach to enhancing deep models’ performance, we employ the mixup method  [20] to expand our training dataset to 5000 samples. This process involves mixing pairs of data and label samples from the original dataset based on weights derived from a symmetric Beta distribution, specifically Beta($\alpha=0.2$, $\beta=0.2$). This augmentation strategy is critical before employing deep generative models, which require substantial amounts of data. By increasing our training dataset size through mixup, we address the high data demands of these models and promote label smoothing. This preparation helps mitigate overfitting and enhances the model’s generalization capabilities by exposing it to a more varied set of training examples, which is essential for effectively training deep generative models to learn robust features. Note that this augmentation strategy is specifically tailored to improve the training of deep generative models and is not directly applicable to classification tasks due to the generation of smooth labels that may obscure distinct class boundaries. CGAN is a type of GAN that conditions its generation process on additional information like labels or data from other modalities. This conditioning leads to more controlled and targeted generation of data, allowing the network to produce specific types of outputs based on the given conditions [31]. We enhance the implementation of CGAN by integrating Wasserstein GAN with Gradient Penalty (WGAN-GP), which improves training stability and mitigates mode collapse, thereby ensuring higher quality and diversity in the generated outputs  [32]. In our architecture of CGAN, the selection of depth and width for layers follows a pyramidical structure, with an increasing or decreasing format to control processing of information, where each layer refines the output of its predecessor. We also employ label injection across all linear layers, not just the initial ones; see Table 3 where $C\in\{4,10\}$ denotes class size. This approach strengthens the labels’ influence throughout the network. This integration of labels into the network’s layers, as evidenced in literature, bolsters the conditioning effect, aligning outputs more closely with specified labels [33, 34, 35]. Our hyperparameter configuration includes the use of the LeakyReLU activation function, set with a slope of 0.2 for negative input values, to prevent the “dying ReLU” problem as well as to capture non-linearities in the data  [36, 37]. Batch size is determined as 64 and the learning rate for the Adam optimizer is set to 0.0005 for both discriminator and generator.
@@ 7 Tanh - - - - - -
CVAE, a variant of VAE, distinctively incorporates conditional data such as labels or contextual inputs to steer its generative mechanism. This incorporation of conditional information is central to CVAE’s functionality, enabling it to generate data that is both precise and contextually relevant, allowing the network to create outputs finely tuned to the specified conditions. Similar to CGAN, our CVAE architecture embodies a tiered layer structure, mirroring a pyramid-like progression in the depth and width of layers with labels incorporated in all linear layers; see Table 4. CVAE utilizes a dual-component loss function: a reconstruction loss, ensuring the fidelity of the generated data to the original input, and a Kullback-Leibler (KL) divergence loss, which maintains the efficiency and effectiveness of the latent space representation  [38, 19]. To dynamically balance the contribution of the KL divergence loss across training epochs, the CVAE employs a cyclical annealing schedule, with the KL weight peaking at its maximum value of 1.0 every 10 epochs. This sinusoidal modulation facilitates a smoother and more controlled integration of the KL divergence into the overall loss, enhancing the stability and quality of the latent space learning process  [39]. The configurations for our model include using a LeakyReLU activation function with a slope of 0.2, a batch size of 64, and an Adam optimizer with a learning rate of 0.0001.
@@ 4.3 Classifier
We opt for a pipeline that integrates the Synthetic Minority Over-sampling Technique (SMOTE)  [40, 41] with XGBoost’s Classifier  [42, 43]. This strategic choice targets class imbalance, a pervasive challenge in machine learning, using SMOTE to bolster the representation of minority classes effectively. Coupled with XGBoost, renowned for its efficiency, scalability, and adaptability across various applications, we employ an ensemble method that significantly enhances accuracy and generalization. This pipeline aligns seamlessly with our overarching research priorities, allowing to channel our efforts and resources more effectively into the core aspect of our work - synthetic data generation - while still ensuring reliable classification results. In our setup, 20% of the training data is reserved for five random validation sets, over which results are averaged to enhance robustness and reliability for every case. We incorporate a patience of 10 evaluations during validation to determine early stopping, further refining our model’s performance. Our model is configured with the “multi:softprob” objective, suitable for multiclass problems as it outputs a probability for each class. It evaluates using the “mlogloss” and “merror” metrics; “mlogloss” measures the accuracy of the probability outputs while “merror” tracks the misclassification rate, together providing a comprehensive view of model performance across multiple classes.
@@ 4.4 Evaluation metrics
The evaluation is conducted on ten tasks: 4-classes and 10-classes UAV (and flight mode) identification, each having five cases of generation on 10, 20, 30, 40, and 50% of the complete data. For each case, we use a randomly sampled 5-fold cross-validation [44] setting and report the average scores over all of the folds. For the evaluation of multilabel classification in every scenario, the accuracy, precision, recall, and F1-score are computed for each label and subsequently averaged based on the support of each label (the number of true instances for each label).
@@ 5.1 Validation of bounded divergence
In Section 3, we established an upper bound for the distributional distance between synthetic and real data, focusing on the distribution of patches within samples. Specifically, the bound presented in Corollary 2 ensures that two-dimensional arrays, such as images, generated by one-shot generative models like GPDM, do not exhibit a distributional divergence from real arrays beyond a predefined threshold. In this subsection, we demonstrate that such threshold holds in practice. Let $W$ represent the distributional distance between synthetic and real datasets, and $\delta$ denote the expected distributional distance between their respective patches. We demonstrate the relationship of $W/\delta$ to its upper bound, as detailed in Eq. (22). Fig. 5 confirms that Eq. (22) is applicable across various scenarios, encompassing different class counts, patch sizes used in generation, proportions of real data utilized, and the patch size $(n^{\prime})$ employed for calculating $\delta$. The figure presents the mean of $W/\delta$, calculated over five folds, and also shows this mean value plus ten times the standard deviation of $W/\delta$. Despite the consistent applicability of the bound across all scenarios, it may not always appear tight. In this work, the observed phenomenon can primarily be attributed to the multiscale approach of GPDM. This approach initially concentrates on capturing the global structure of samples at a lower scale. This process effectively reduces $W$ to a greater extent. Also, $\delta$ is subject to overestimation using limited patches within each sample, because the Wasserstein distance is sensitive to the tails of distributions. Limited samples may not adequately represent these tails, leading to an overestimation of the actual distributional distance. Additionally, in high-dimensional spaces, the curse of dimensionality can inflate perceived differences, further contributing to overestimation. Sampling variability, especially when capturing more extreme values or outliers, can also skew the perceived distribution, resulting in an overestimated Wasserstein distance between the respective patches $(\delta)$. Therefore, we have successfully established and validated an upper bound for the distributional distance between synthetic and real data using one-shot generative models, specifically focusing on GPDM. Our investigations confirm the practical applicability of the established bound across a range of scenarios, as demonstrated in the consistency of $W/\delta$’s adherence to its theoretical upper limit. However, it is important to recognize that certain factors, such as the multiscale nature of GPDM and the limited patch counts within each sample, can lead to the underestimation of $W/\delta$’s in empirical circumstance.
@@ 5.2 Classification performance
We can evaluate the identification of UAV types and flight modes by employing synthetic RF transforms across various scenarios. These scenarios are distinguished by the proportion of data used in the synthesis process and the generation methodology. Specifically, we assess classification performance by comparing a complete dataset and its subsets sampled at different proportions with their corresponding synthetically generated counterparts, utilizing a range of models. The classification performance metrics are presented in Figs. 6 and 7, which both incorporate results from complete and subsampled datasets for comparison. In this context, the complete datasets serve as our ground truth, while the subsampled datasets function as baselines. The metrics for the complete data remain constant since data proportion is not a factor. On the other hand, the metrics for the subsampled data improve with the use of more data.
@@
Figure 6: The performance of UAV identification using synthetic transformed RF signals generated by GPDM.
@@
Figure 7: The performance of UAV identification using synthetic transformed RF signals generated by CGAN and CVAE.
@@
Fig. 6 illustrates the performance metrics of classification using synthetic RF transformations generated by the one-shot GPDM approach. Generally, the performance metrics of the synthetic data fall between those of the baseline and the ground truth, demonstrating the effectiveness of the one-shot generative approach for data synthesis. The impact of patch size on various metrics is significant, especially in the 10-class scenario where utilizing a patch size of 5 is more successful than larger patch sizes in expanding the support of class-wise distributions. The benefit of choosing a smaller patch size becomes less pronounced as the proportion of data accessed increases (approaching 50%), where further extrapolation of distributional support no longer offers additional benefits. Fig. 7 illustrates the performance metrics of classifications using synthetic RF transformations generated by conditional deep learning models such as CGAN and CVAE. Contrary to the GPDM approach, CGAN and CVAE underperform, often not even reaching the baseline metrics, primarily due to their inefficiency in scenarios with limited data. In such low-data regimes, the outcome where deep learning models tend to memorize small training samples represents the best-case scenario. This already precarious situation is further aggravated if the training data is not meticulously clean. The presence of noise and outliers in the data poses significant challenges, making it a struggle for these models to even achieve basic memorization of the training samples, let alone generalize effectively. One-shot generative data augmentation, in contrast to CGAN and CVAE, excels particularly in environments with outliers. Its robustness arises from its capacity for internal learning, where the model extracts and utilizes the intrinsic characteristics of a single example to generate a broad range of synthetic samples. This approach focuses on the core attributes of the sample, effectively minimizing the influence of outliers. By capturing and extrapolating the essential features, the one-shot method not only improves the diversity and representativeness of the synthetic data but also significantly enhances its utility for training models in scenarios plagued by data anomalies. The comparative analysis, depicted in Fig. 6 and Fig. 7, reveals significant insights into the performance of the one-shot generative approach using GPDM and conditional deep generators such as CGAN and CVAE. While the one-shot GPDM method consistently demonstrates superior handling and generation of synthetic data in the presence of outliers and limited data scenarios, the performance of CGAN and CVAE lags, especially when data cleanliness and abundance are issues. This study underscores the critical importance of choosing appropriate data synthesis techniques based on the specific challenges and data environment encountered. It suggests that one-shot generative methods, with their robust internal learning mechanisms, are not only more resilient but also more adaptable to diverse and challenging data landscapes, making them a preferable choice in situations where data quality cannot be guaranteed.
6 Limitations and Future Works
The primary focus of our research was to address data scarcity in RF-based UAV identification through one-shot generative data augmentation, but we did not analyze how individual environmental factors affect performance. Future work could involve a detailed analysis of the model’s performance against datasets that incorporate factors such as multipath propagation, non-line-of-sight conditions, heavy urban interference, or severe weather individually, providing a more comprehensive evaluation of the method’s robustness. Additionally, incorporating a physics-based approach could enhance the generation of more realistic samples, allowing for an insightful understanding of the RF environment’s impact on signal integrity. While one-shot generative data augmentation shows promise in low-data regimes, its computational intensity could pose challenges in specific operational contexts. This issue stems from the inherent complexity of one-shot generative models, suggesting that future research should focus on optimizing these algorithms and exploring high-performance computing techniques. Such advancements could enhance the model’s applicability to real-time identification tasks, also facilitating more extensive studies using datasets that capture varied real-world conditions, thereby further validating its robustness and practical applicability. Lastly, the ethical considerations in deploying one-shot generative data augmentation, particularly in UAV identification, also warrant careful attention. Ensuring that synthetic data does not inadvertently introduce bias or inaccuracies that could compromise UAV detection or misclassify benign entities as threats is paramount. Furthermore, as these models may be employed in security-sensitive areas, the implications of errors—whether false positives or negatives—could have serious repercussions. Therefore, it is essential to establish rigorous validation frameworks and transparency in model training processes to maintain trust and reliability in the use of synthetic data in UAV identification systems.